After you analyze the loop in the following program, find the output.

int y = 3;

for (int i = 1; i < 7; ++i)

{

  y += i;

}

cout << "The value for y is "<<y ;
Answer #1

Initially y is 3.

Let us analyse for loop,

The statement in the loop is y+=i. It means y=y+i.

  • i is 1 and 1 is less than 7. So, y=3+1 = 4. Now y is 4. And i is pre-incremented. So, i is increased to 2 and i is 2.
  • i is 2 and condition is true. So y=4+2=6. And i is now 3.
  • Again 3 lessthan 7. So condition is true. So, y=6+3=9. Now i is 4.
  • Again 4 lessthan 7. So, y=9+4=13. I is 5.
  • Again y=13+5=18. Now i is 6.
  • Again y=18+6=24. Now is 7.
  • In thia case i is 7. But 7<7 is false. As condition is false, we exit from for loop.
  • So, now after coming out of for loop, y value is 24. Here, we can observe that for loop executes for 6 times.
  • Output is displayed as The value for y is 24.

Upvote if u like it.

Thank you.